OpenStudy (anonymous):

Convert 22/125 to a percent

OpenStudy (anonymous):

would it be

OpenStudy (anonymous):

22/125=0.176=176/1000=17.6/100=17.6%

OpenStudy (stormfire1):

\[\frac{22}{125}=0.176\]To convert that to a percent you just multiply it times 100 to get 17.6%.

OpenStudy (stormfire1):

Your method is also correct but took an extra step.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it would be 100*0.176

OpenStudy (stormfire1):

Yes...you get the same answer but it's a little shorter to do.

OpenStudy (anonymous):

\[\frac{22 \times 100 }{125} = 17.6\]

OpenStudy (stormfire1):

which is the same as \[\frac{22}{125}*100=17.6%\]

OpenStudy (anonymous):

which is also the same as \[100 \times \frac{22}{125} = 17.6\]